cos? o + 4 cos a +1=0 interval (0° 3000) I X=?
Answer #1

Given i O E[0, 360) cos²o + 4 cos o t =0 x =0=2 we can i write cos + 4 coso + 1 = 0 cose + q los ot 4 -3 = 0 cose+ (2) ²+ 2.(cos 6+ 2)2 = 3 ( Cos@+2) = 3 coso +2 = + 3 coso – 2 & 3 coso = -2+√3 -2-√3 coso Cannot be bess them / a Coso = √3-2 coso = √For mal O = 2xix I 105,5 0=465.5 7 360 O = 2 r I losos o = 360 I losos o=360-10505 0 = 25405 Hemce Answer. Oa 105.5 254.5 L